OpenStudy (mathmath333):

A speaks the truth 3 times out of 4, B 7 times out of 10. They both assert that a white ball has been drawn from a bag containing balls of 6 different colors. Find is the probability of the truth of the assertion?

OpenStudy (studygurl14):

What's your question?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ A speaks the truth 3 times out of 4, B 7 times out of 10.}\hspace{.33em}\\~\\ & \normalsize \text{ They both assert that a white ball has been drawn from a }\hspace{.33em}\\~\\ & \normalsize \text{ bag containing balls of 6 different colors. Find is the probability }\hspace{.33em}\\~\\ & \normalsize \text{ of the truth of the assertion?}\hspace{.33em}\\~\\ & a.)\ \dfrac{12}{49} \hspace{.33em}\\~\\ & b.)\ \dfrac{3}{10} \hspace{.33em}\\~\\ & c.)\ \dfrac{21}{40} \hspace{.33em}\\~\\ & d.)\ \normalsize \text{None of these} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (desmarie):

if you multiply across it would be c

OpenStudy (studygurl14):

It's not that simple @desmarie

OpenStudy (mathmath333):

@Jaydid

OpenStudy (mathmath333):

both A and B are mutually exclusive ?

OpenStudy (mathmath333):

r us sure it is d.)

OpenStudy (mathmath333):

@jayzdd